# Grid Monitoring and Protection with Continuous Point-on-Wave Measurements and Generative AI

### v1

@@ Abstract
Purpose: This article presents a case for a next-generation grid monitoring and control system, leveraging recent advances in generative artificial intelligence (AI), machine learning, and statistical inference. Advancing beyond earlier generations of wide-area monitoring systems built upon supervisory control and data acquisition (SCADA) and synchrophasor technologies, we argue for a monitoring and control framework based on the streaming of continuous point-on-wave (CPOW) measurements with AI-powered data compression and fault detection.
Methods and Results: The architecture of the proposed design originates from the Wiener-Kallianpur innovation representation of a random process that transforms causally a stationary random process into an innovation sequence with independent and identically distributed random variables. This work presents a generative AI approach that (i) learns an innovation autoencoder that extracts innovation sequence from CPOW time series, (ii) compresses the CPOW streaming data with innovation autoencoder and subband coding, and (iii) detects unknown faults and novel trends via nonparametric sequential hypothesis testing.
Conclusion: This work argues that conventional monitoring using SCADA and phasor measurement unit (PMU) technologies is ill-suited for a future grid with deep penetration of inverter-based renewable generations and distributed energy resources. A monitoring system based on CPOW data streaming and AI data analytics should be the basic building blocks for situational awareness of a highly dynamic future grid.
@@
Grid monitoring, Generative AI, Continuous-point-on-wave measurements, Synchro-waveform, Power signal compression, Fault detection, and Overcurrent Protection.
@@ 1 Introduction
Over the past decades, the power grid has undergone unprecedented changes, from one with high-inertial synchronous generators dispatched to follow predictable demands to one with substantial non-dispatchable stochastic renewable resources integrated into the grid via power electronics operating at the nano-second time scale. With highly variable inverter-based resources (IBRs) such as wind/solar and reduced system inertia, the grid is increasingly stochastic and dynamic, frequently operating near its limits. Externally, accelerated occurrences of climate-related events put enormous stress on the aging grid. Outage events increased by 78% in the decade of 2011-2021 over the previous one [1], causing an average annual economic loss of over 150 billion dollars [2].
Not all outages can be prevented, but it begs the question: How many outages can be mitigated, load sheds restored quickly, and economic impacts minimized with today’s technology and advances in computing, networking, and artificial intelligence?
This paper makes a case for a next-generation AI-powered grid monitoring system based on continuous point-on-wave (CPOW) measurements, providing high-resolution and high-fidelity grid monitoring and protection beyond the capability of the state-of-the-art synchrophasor technology. To this end, we present an AI-enabled data-streaming compression architecture and data-driven monitoring algorithms to detect anomalies and novel trends. Our approach is inspired by the classic Wiener-Kallianpur innovation representation of random processes [3, 4] and recent advances in generative AI capable of producing artificial random samples matching the originals.
1.1 Beyond PMU with CPOW Measurements
At the outset, it is highly relevant to address a cogent question: why should one consider new types of measurements when the phasor measurement units (PMUs) and synchrophasor technology can already provide synchronized measurements with resolutions higher than the state-of-the-art SCADA measurements? What can CPOW measurements reveal that SCADA and PMU measurements cannot?
PMU was invented over thirty years ago as a technology for wide-area situational awareness. The number of PMUs deployed in utilities in the US finally surpassed 3000 in 2022, far from adequate to cover the operation footprint of a small to medium-sized system operator. After three decades of (mostly governmental) investments worldwide and decades of extensive research, the technology has yet to take off. The unfulfilled promise of PMU is not because its technology is ahead of its time; PMU is based on the digital signal processing and the global positioning technologies of the seventies and eighties; both are inexpensive and widely used in commercial sensing and communication devices, including miniature wearable devices such as smartwatches and earbuds. It is difficult to foresee that PMU technologies can make significant impacts on reducing and preventing future blackouts.
@@ 1.1 Beyond PMU with CPOW Measurements
One technical reason for the lack of deployment is that the 10 to 250 Hz reporting rates of the state-of-the-art PMU cover only a small fraction of significant grid events, as shown in Fig. 1. In particular, PMU extracts current and voltage phasors associated with the fundamental frequency of 50 or 60 Hz, removing the frequency contents outside a narrow frequency band around the fundamental frequency. As such, PMU measurements are insensitive to most grid event waveforms shown in Fig. 1, unable to capture critical grid events when the network is under stress while producing too much data when the grid is normal. There is also the economic reason for the chicken-and-egg trap: insufficient monitoring resolution and coverage make PMU ineffective as a practical wide-area monitoring of impactful events, and the lack of compelling use cases limits necessary investments and broad deployments to realize its potential.
Interestingly, CPOW data are ubiquitous in modern grids in the form of direct Nyquist-sampled current and voltage waves. Measurements at all digital sensors, intelligent electronic devices (IEDs), PMUs, digital relays, and digital fault recorders are all CPOW data with sufficiently high sampling rates to capture highly dynamic signal waveforms from grid events. These high-resolution data rarely go beyond substations for grid situational awareness. Why not use CPOW data for high-resolution monitoring and control then? The common wisdom has been that (i) streaming CPOW measurements requires substantial infrastructure investments and bandwidth allocation; (ii) when the grid is in normal operating conditions, even PMU measurements are not ”useful” most of the time. Why go with a technology with 10 to 100 times higher data rates?
The inadequate bandwidth argument against high-resolution CPOW streaming is unconvincing with today’s technology. Since the invention of PMU, the rate of communications has increased by nearly six orders of magnitude, and the world of communications has grown from zero cell phones to 5G/6G digital cellular, from text-based communications to real-time streaming of live sports events. Underlying all these are signal processing and networking technologies that deliver high-fidelity entertainment and real-time control of robots, autonomous vehicles, and drones. If there are data communications technology barriers, they are by no means insurmountable.
The economic argument deserves more careful deliberation. Indeed, it seems wasteful to allocate bandwidth for contingencies that occur on rare occasions. It is also true that efficient data analytic methods are needed to sieve through large amounts of data in searching for early signs of network instability and actionable information. However, not having high-resolution data essential for a broad spectrum of grid events inhibits advances that can lead to breakthrough technologies critical to grid reliability and resiliency.
While not deployed in practice, the idea of using CPOW data for grid monitoring has been considered in the literature in recent years [9, 10, 11, 6, 12]. Recent reviews of synchronized CPOW technology (a.k.a synchro-wave) are [13, 14], where the authors articulate broad visions of using CPOW data for monitoring, protection, and control. The work presented here aligns with these visions. We provide a possible architectural design and address data analytic challenges when measurements are highly stochastic.
@@ 1.2 Scope, Contributions, and Organization
This paper addresses two challenges in future grid monitoring systems: (i) efficient compression and communications for real-time streaming of CPOW measurements within the current bandwidth limits for PMU streaming technology and (ii) rapid and reliable detection of faults and novel trends from highly stochastic CPOW measurements.
The main contribution of this work is twofold. First, we develop a novel compression architecture that combines machine-learning-based data analytics and data compression for CPOW data streaming. Such technology is critical for bandwidth-efficient, dynamic, and interactive operations among IEDs, substations, and the control center during severe grid events. Second, we propose a sequential detection approach for the reliable and rapid detection of anomalies and novel trends from CPOW data streaming. Built on the classic innovation representation of time series pioneered by Wiener and Kallianpur, the proposed generative machine learning technique distinguishes the normal and the faulty in the feature domain of the underlying probability distributions of CPOW data, in contrast to conventional methods that directly compare current and voltage phasors with their nominal values. Finally, we apply the proposed anomaly detection approach to over-current protection in distribution systems, where the presence of distributed stochastic generations causes protection blinding and sympathetic tripping. We demonstrate 16% improvements in detection accuracy and 67.9% in detection speed.
This paper is organized as follows. Section 2 describes a generative AI approach to representation learning of high-resolution time series (such as that from CPOW data streaming), which transforms the generic nonparametric time series to a canonical independent and identically distributed (IID) innovation sequence by an innovation autoencoder. This particular autoencoder structure plays a key role in both CPOW data compression and anomaly detection. In particular, anomaly detection problems are formulated and solved as nonparametric hypothesis testing with arbitrary and unknown distributions. Section 2.3 presents a machine-learning (ML) based data compression aimed at achieving the optimal compression rate-distortion tradeoff. Section 3 presents the application of the proposed anomaly detection in over-current protection.
@@
Notations used in this work are standard. Random variables are nominally denoted by capital letters, and their realizations in small letters. Vectors, matrices, and sequences are usually in boldface. We use ${\bm{X}}=(X_{t})$ to represent a random sequence and ${\bm{x}}=(x_{t})$ its realization. When considering a segment of a random sequence, ${\bm{X}}_{t_{1}:t_{2}}$ denotes the segment from $t_{1}$ to $t_{2}$, i.e., ${\bm{X}}_{t_{1}:t_{2}}=(X_{t_{1}},\cdots,X_{t_{2}})$. For two random variables $X$ and $Y$, $X\stackrel{{\scriptstyle\mbox{\sf\small a.s.}}}{{=}}Y$ means the two equal almost surely and $X\stackrel{{\scriptstyle\mbox{\sf\small d}}}{{=}}Y$ equal in distribution. Table 1 gives a list of designated symbols.
Table 1: Mathematical notations used in this paper.
$(X_{t})$ The random process representing a sequence of CPOW measurements.
$(V_{t})$ The innovation sequence.
$(U_{t})$ An IID sequence of uniform distribution.
$(\hat{X}_{t})$ The rescontructed sequence output by IAE decoder.
$(x_{t})$ A sequence of real numbers indicating the past observations of $({\bm{X}}_{t})$.
$G$ Innovation encoder function.
$H$ Innovation decoder function.
$G_{\theta}$ A neural network approximation of $G$ parameterized by $\theta$.
$H_{\eta}$ A neural network approximation of $H$ parameterized by $\eta$.
$D_{\gamma}$ Innovation discriminator that measures the distance between $({\bm{V}}_{t})$ and $({\bm{U}}_{t})$.
${\cal U}[0,1]$ The continuous univariate uniform distribution on $[0,1]$.
2 Learning Innovation Representations
The proposed AI-powered monitoring and anomaly/novelty detection approaches are rooted in the classic idea of innovation representation of time series first proposed by Wiener and Kallianpur [3], who made a profound conjecture that stationary random processes can be represented by an IID uniform (or Gaussian) innovation sequence. By removing temporal dependencies and converting a generic stationary process to a canonical IID sequence with simple uniform/Gaussian marginals, the innovation representation is arguably the most efficient representation of a nonparametric random process.
2.1 Wiener-Kallianpur Innovation Representation
In the parlance of modern machine learning, an innovation representation of a stationary random process $(X_{t})$ is a causal autoencoder shown in Fig. 2, with the latent process $(V_{t})$ being an IID-uniform innovation sequence. As suggested by the word innovation, $V_{t}$ represents the new information contained in $X_{t}$ not present in (i.e., statistically independent of) the past ${\bm{X}}_{0:t}=\{X_{t-1},X_{t-2},\cdots\}$.
Figure 2: An autoencoder interpretation of innovation representations.
Mathematically, the innovation representation of the time series is defined by a pair of causal mappings $(G,H)$ and the innovation sequence $(V_{t})$:

### v2

# Grid Monitoring with Synchro-Waveform and AI Foundation Model Technologies

@@ Abstract
Purpose: This article advocates for the development of a next-generation grid monitoring and control system designed for future grids dominated by inverter-based resources. Leveraging recent progress in generative artificial intelligence (AI), machine learning, and networking technology, we develop a physics-based AI foundation model with high-resolution synchro-waveform measurement technology to enhance grid resilience and reduce economic losses from outages.
Methods and Results: The proposed framework adopts the AI Foundation Model paradigm, where a generative and pre-trained (GPT) foundation model extracts physical features from power system measurements, enabling adaptation to a wide range of grid operation tasks. Replacing the large language models used in popular AI foundation models, this approach is based on the Wiener-Kallianpur-Rosenblatt innovation model for power system time series, trained to capture the physical laws of power flows and sinusoidal characteristics of grid measurements. The pre-trained foundation model causally extracts sufficient statistics from grid measurement time series for various downstream applications, including anomaly detection, over-current protection, probabilistic forecasting, and data compression for streaming synchro-waveform data. Numerical simulations using field-collected data demonstrate significantly improved fault detection accuracy and detection speed.
Conclusion: The future grid will be rich in inverter-based resources, making it highly dynamic, stochastic, and low inertia. This work underscores the limitations of existing Supervisory-Control-and-Data-Acquisition and Phasor-Measurement-Unit monitoring systems and advocates for AI-enabled monitoring and control with high-resolution synchro-waveform technology to provide accurate situational awareness, rapid response to faults, and robust network protection.
@@
Generative AI, AI Foundation Model, Continuous-point-on-wave measurements, Synchro-waveform measurements, Power system monitoring and control, Power signal compression, Fault detection, and Over-current Protection.
@@ 1 Introduction
Over the past decades, the power grid has undergone unprecedented changes, evolving from a system dominated by high-inertia synchronous generators dispatched to meet predictable demands to one with substantial non-dispatchable, stochastic, and inverter-based renewable resources (IBRs) such as wind, solar, and battery storage devices. Operating at microsecond timescales, these resources reduce system inertia, making the grid more stochastic and dynamic, frequently operating near its limits.
Externally, the accelerated climate-related events place enormous stress on the aging grid. Outage events between 2011 and 2021 increased by 78% compared to the previous decade [1], resulting in an average annual economic loss exceeding 150 billion dollars [2]. While not all outages can be prevented, this raises a critical question: how many outages can be mitigated, load sheds restored quickly, and economic impacts minimized by advanced computing, networking, and artificial intelligence (AI)?
This paper advocates for a next-generation AI-powered grid monitoring system built on synchro-waveform measurement technology, enabling high-resolution and high-fidelity monitoring beyond the capabilities of state-of-the-art SCADA [^1] Supervisory Control and Data Acquisition and Phasor Measurement Unit (PMU) technologies. To achieve this, we introduce a physics-based AI Foundation Model framework, augmenting the popular Large Language Model (LLM) used in most AI foundation models for natural language processing with the classic Wiener-Kallianpur-Rosenblatt Innovation Representation Model (IRM) [3, 4] for power system signal processing for monitoring, control, and real-time decision-making.
1.1 From Synchro-phasor to Synchro-waveform Technology
The synchro-waveform technology generates direct and synchronously sampled current and voltage waveform measurements, and a device that produces such direct measurements is called Waveform Measurement Units (WMU) [5]. WMUs are ubiquitous in modern grids. Digital sensors, intelligent electronic devices (IEDs), PMUs, digital relays, and digital fault recorders all inherently generate waveform-sampled data with sufficiently fast sampling rates to capture the highly dynamic waveforms of grid events before converting them to statistics (such as phasors) for monitoring tasks. If the next-generation monitoring system is based on high-resolution synchro-waveform measurements, compelling arguments are needed for adopting new WMU in favor of the mature PMU and SCADA technologies. What can WMU measurements reveal that SCADA and PMU measurements cannot? Do the benefits of WMU data justify the costs of new devices and infrastructure investments?
The PMU technology was introduced over three decades ago as a transformative successor to the SCADA technology for wide-area situational awareness in power systems. Despite its potential, PMU deployment in U.S. utilities remains limited, with only a few thousand units installed to date—insufficient for high-resolution monitoring across the operational footprint of even small to medium-sized system operators. After decades of substantial global investment, primarily driven by governmental initiatives and extensive academic and industrial research, the technology has yet to achieve its anticipated impact. In light of this stagnation, it is difficult to envision PMU technology making a transformative impact in addressing emerging challenges in an IBR-dominated power grid.
@@ 1.1 From Synchro-phasor to Synchro-waveform Technology
The barrier to entry for the PMU technology is twofold. As a technology, PMU has a limited reporting rate, capturing only a small fraction of significant grid events, as illustrated in Fig.1. A PMU extracts current and voltage phasors associated with the fundamental frequency of 50 or 60 Hz, filtering out frequency components beyond a narrow band around the fundamental frequency. Consequently, PMU measurements are insensitive to high-order and intra-harmonics of many critical grid events shown in Fig.1, inadequate to monitor critical dynamics during network stress while generating excessive data under normal operating conditions. The so-called chicken-and-egg economic barrier also exists: insufficient monitoring resolution and limited network coverage make PMUs ineffective for practical wide-area monitoring of impactful events, which limits the development of compelling use cases. The lack of convincing applications limits investments and deployments necessary to unlock the technology’s full potential.
Two key arguments are often raised against adopting the higher-resolution synchro-waveform technology. The first is the cost of infrastructure investments and bandwidth requirements to stream these data. However, this argument is increasingly unconvincing in light of advancements in compression, communications, and networking technologies over the last two decades. Since the invention of PMU, communication rates have increased by nearly six orders of magnitude. Today’s communication networks can seamlessly handle high-fidelity entertainment, real-time streaming of live events, and control of autonomous systems such as drones and robots. The notion that streaming synchro-waveform data is infeasible is no longer convincing.
The second argument pertains to the perceived benefits of using synchro-waveform data. The PMU technology has been criticized for contributing to a ‘data deluge’ during normal grid operation, and the WMU technology would generate 10 to 100 times more data, which appears to be a considerable waste of resources devoted to relatively rare occasions. A counterargument is that the grid is not always in a normal or failure state; often, the transition from normal to failure involves sequences of incipient faults that are sporadic and sometimes self-clearing. Detecting such faults requires high-resolution measurements. Discounting the potential of high-resolution measurement technology stifles progress toward breakthrough technologies critical for grid reliability and resiliency.
More convincingly, the challenge of managing data deluge lies in the technology that effectively sifts through large datasets to identify early signs of instability and actionable information. A practical solution to balance the need for high-resolution data and avoiding data overload is the hierarchical push-pull networking: push data from WMUs to the control center when local events are detected, pull data from WMUs when global assessment is necessary. The mathematical foundations and practical implementations of such protocols for content delivery networking have long existed in the literature. See, e.g., [10, 11].
While the WMU technology has not been deployed and effective data analytic techniques missing, the benefits of using synchro-waveform have been well articulated for almost a decade, see e.g., [12, 13, 14, 15, 16, 17]. Recent reviews of synchro-waveform measurement technology [17, 5] have articulated broad visions for its use in monitoring, protection, and control. This article aligns with these visions, offering a possible architectural design within the modern AI foundation model paradigm and addressing the data analytic challenges posed by highly stochastic measurements.
@@ 1.2 Scope, Contributions, and Organization
This paper develops key components of an AI-enhanced WMU-based grid monitoring system, including (i) a novel foundation model architecture using the innovation representation model (IRM) for synchro-waveform time series, (ii) the pre-training of the IRM-based foundation model, and (iii) foundation model adaptations for anomaly detection, over-current protection, and synchro-waveform data compression.
The main contributions of this work are as follows.
  1. 1.
    An IRM-based foundation model architecture: A fundamental challenge in compressing and performing statistical inference on synchro-waveform streaming is the unknown probability models of WMU data. To address this, we propose the innovation representation model (IRM) based on the classic work of representation of random processes by Norbert Wiener and Gopinath Kallianpur [3, 4] and Murray Rosenblatt [18]. IRM provides a unified data processing foundation for all statistical decision tasks, including the monitoring, protection, probabilistic forecasting, and compression applications considered here.
  2. 2.
    Rapid detection of faults and novel trends: We introduce a sequential detection method that identifies anomalies and novel trends in WMU data streams. Unlike conventional methods that rely on direct comparisons of current and voltage phasors to nominal values, our approach operates in the innovation feature space. A novel kernel technique based on a sequential adaptation of the classic Neyman Smooth Test (NST) ensures rapid detection with high confidence.
  3. 3.
    Application to over-current protection: The proposed anomaly detection method is applied to over-current protection in distribution systems, addressing critical challenges such as protection blinding and sympathetic tripping caused by stochastic distributed generation. We demonstrate significant improvements in both detection accuracy and speed.
  4. 4.
    Data compression for synchro-waveform streaming: We develop a novel adaptive subband compression architecture that integrates innovation representation for compressing WMU data. By dividing WMU measurements into harmonic and inter-harmonic subbands, the proposed compression approach compresses and communicates selected active subbands based on their levels of innovation features. This method is crucial for enabling bandwidth-efficient, dynamic, and interactive operations among intelligent electronic devices (IEDs), substations, and control centers, especially during severe grid events.
This paper is organized as follows. Sec. 2 introduces an AI foundation model approach to power system monitoring and control based on the innovation representation of power system measurements. We discuss the architecture of the proposed foundation model and foundation model pre-training using WMU data for downstream foundation model adaptations. Three foundation model adaptations are then discussed in subsequent sections. Sec. 3 presents the adaptation of the proposed foundation model for anomaly detection. A novel sequential detection scheme based on Neyman’s smooth test is derived to achieve fast detection speed and high accuracy. Sec. 4 presents the foundation model adaptation for over current protection. Sec. 5 presents the foundation model adaptation for the event-driven compression of WMU measurements.
@@
Notations used in this work are standard in statistical learning and power system literatures. Vectors, matrices, and sequences are usually in boldface. When considering a segment of a sequence ${\bm{x}}$, ${\bm{x}}_{t_{1}:t_{2}}$ denotes the segment from $t_{1}$ to $t_{2}$, i.e., ${\bm{x}}_{t_{1}:t_{2}}=(x_{t_{1}},\cdots,x_{t_{2}})$. When necessary to distinguish a random variable and its realization, a random variable is capitalized and its realization in small letter. For two random variables $X$ and $Y$, $X\stackrel{{\scriptstyle\mbox{\sf\small a.s.}}}{{=}}Y$ means the two equal almost surely and $X\stackrel{{\scriptstyle\mbox{\sf\small d}}}{{=}}Y$ equal in distribution. Table 1 gives a list of designated symbols and key abbreviations.
Table 1: Major Nomenclatures and Mathematical Symbols.
DG/SDG Distributed generation/Stochastic distributed generation.
IBR Inverter-based resources.
IID: Independent and Identically Distributed
IID-Uniform IID random variables with uniformly distributed marginals
LLM/IRM: Large Language Model/Innovation Representation Model.
NST/SNST: Neyman’s Smooth Test/Sequential Neyman’s Smooth Test.
PMU/WMU: Phasor Measurement Unit/Waveform Measurement Unit.
$(x_{t})$: a sequence of WMU measurements.
$(\nu_{t})$: an innovation sequence.
$(u_{t})$: the IID sequence with the uniformly distributed marginal.
$(\hat{x}_{t})$: the reconstructed $(x_{t})$ at the output by the foundation model autoencoder.
$G$: the innovation encoder function.
$H$: the innovation decoder function.
$G_{\theta}$: the neural network representation of $G$ parameterized by $\theta$.
$H_{\eta}$: the neural network representation of $H$ parameterized by $\eta$.
$D_{\gamma}$: the neural network representation of innovation discriminator between $(\nu_{t})$ and $(u_{t})$.
${\cal U}[0,1]$: the continuous univariate uniform distribution on $[0,1]$.
2 A Foundation Model for Synchro-waveform Data
AI Foundation Model, broadly defined, is a learning and decision architecture that is trained on vast data and adaptable to a wide range of applications [19]. Some of the defining features of foundation model include the following:
  • Architectural partition of pre-training and the adaptation: The foundation model approach to AI partitions the complex and costly process of foundation model pre-training—training deep neural networks with billions of parameters and vast training data—from foundation model adaptations through fine-tuning for specific applications. Such “division of labor” is essential to market general AI technologies to a broad range of applications.
  • Self-supervised learning of essential features: For a foundation model to be effective for a wide range of applications means that the foundation model must capture essential features in vast data, making self-supervised learning essential with unlabeled training data.
  • Generative AI: Foundation models are generative, capable of producing artificial samples not in the training data but with the same characteristics. The generative capability of a foundation model is crucial for generalization and data augmentation.
Most successful AI foundation models are based on large language models (LLMs), pre-trained with vast text, image, and video data. The transformer architecture and attention mechanisms play a critical role in these foundation models to capture naturally highly complex linguistic structures without relying on grammatical rules or stylistic constraints. A key feature of LLM is its ability to model long-range (passage) linguistic dependencies in language.
The adoption of the AI foundation model principle to power systems is well recognized and emerging but still limited. See, e.g., [20, 21, 22] and references therein. This work is the first attempt to develop an AI foundation model for synchro-waveform data generated by WMUs. Key challenges for data produced by the grid arise from the fundamental differences between grid-generated data and language texts; the former follows the physical laws of Ohms, Kirchhoff, and Maxwell, and the latter the linguistic structures and conventions. The power system signals are sinusoidal; they are generated causally through a physical process. Typically, they do not have the long-range dependencies as in linguistic passages. It is not clear that the principle of LLM-based foundation model is suitable to power system applications involving measurements of physical (rather than linquistic) attributes.
2.1 Innovation Representation Model
A generic foundation model structure is an autoencoder that maps the foundation model input to a latent (feature) process by the encoder, and the foundation model input is approximated by the decoder to ensure that the latent feature is, in some sense, a faithful representation of the input. Typically, the latent process is simpler and has lower dimensionality. For the LLM-based foundation model, the autoencoders are various types of transformers with attention mechanisms.
We propose an innovation-based foundation model for time series from power systems for monitoring and control, having all the general characteristics of foundation model but augmenting the linguistic-based LLM autoencoder with a time series-based Innovation Representation Model (IRM), denoted by $({\cal G}_{\theta},{\cal H}_{\eta})$, as shown in Fig. 2. While LMM-based foundation may be important for broad power system applications, here we focus on power system monitoring and real-time decision-making using power system measurements, where the porposed IRM-based foundation model can be more effective.
Figure 2: The foundation model pre-training and adaptations with innovation representation model (IRM).